# Learning production functions for supply chains with graph neural networks

### v1

@@ Abstract
The global economy relies on the flow of goods over supply chain networks, with nodes as firms and edges as transactions between firms. While we may observe these external transactions, they are governed by unseen production functions, which determine how firms internally transform the input products they receive into output products that they sell. In this setting, it can be extremely valuable to infer these production functions, to better understand and improve supply chains, and to forecast future transactions more accurately. However, existing graph neural networks (GNNs) cannot capture these hidden relationships between nodes’ inputs and outputs. Here, we introduce a new class of models for this setting, by combining temporal GNNs with a novel inventory module, which learns production functions via attention weights and a special loss function. We evaluate our models extensively on real supply chains data, along with data generated from our new open-source simulator, SupplySim. Our models successfully infer production functions, with a 6-50% improvement over baselines, and forecast future transactions on real and synthetic data, outperforming baselines by 11-62%.
@@ 1 Introduction
Supply chains form the backbone of the global economy and disruptions can have enormous consequences, costing trillions of dollars (Baumgartner et al.,, 2020) and risking national security (The White House,, 2021). Thus, modeling supply chains and predicting how they will evolve, especially under shocks, are essential. To model supply chains, we can represent them as temporal (or dynamic) graphs, where nodes are firms and edges represent time-varying transactions between firms. A unique feature of these graphs is that they are governed by underlying production functions: firms receive input products (e.g., wheels) from other firms, internally transform those inputs into outputs (e.g., cars) via production functions, then sell those output products to other firms or consumers (Carvalho and Tahbaz-Salehi,, 2019). These production functions dictate which firms are connected to each other, as well as the timing of transactions across the network. For example, if a firm experiences a shortage in an input, its production of outputs that rely on that input will be disrupted, but it can continue to produce other outputs for some time. Conversely, if a firm needs to produce more of a particular output, due to a change in demand, it will increase its orders of the inputs required to produce that output, but orders for other inputs may remain unchanged.
@@
Despite growing interest in developing graph neural networks (GNNs) for temporal graphs (Huang et al.,, 2023; Longa et al.,, 2023), temporal graphs governed by production functions—which we term temporal production graphs (TPGs) in this work—have not yet been explored. Beyond supply chains, examples also appear in biological domains, such as enzymes producing metabolites within metabolic pathways, or in organizations, where teams rely on inputs from other teams to produce their outputs (e.g., engineers interfacing with product designers). TPGs introduce new challenges compared to other temporal graphs, such as the need to infer the production functions. Inferring a TPG’s production functions can help us better understand the graph’s mechanisms and intervene more effectively, such as identifying bottlenecks in supply chains and proposing plausible solutions to improve efficiency and sustainability (Aigner and Chu,, 1968; Coelli et al.,, 2005). Furthermore, even standard tasks, like link prediction, may be complicated under TPGs. For example, while standard temporal GNNs might capture disruptions generally propagating across connected firms in supply chains, they will miss the specific connections between each firm’s inputs and outputs and, under a shortage of inputs, not be able to precisely predict which outputs will be affected.
@@
To fill this gap, here we introduce a new class of GNNs designed for TPGs, focusing on two objectives: (1) learning the graph’s internal production functions, (2) predicting its future edges. Prior temporal GNNs have focused primarily on the second objective, but are not designed for the first. Our models support both objectives, by combining temporal GNNs with a novel inventory module, which learns production functions by explicitly representing each firm’s time-varying inventory and updating it based on attention weights that map external supply to internal consumption. Our module can be combined with any GNN; to demonstrate this, here we combine it with two popular temporal GNNs, Temporal Graph Network (TGN) (Rossi et al.,, 2020) and GraphMixer (Cong et al.,, 2023), which we also extend in important ways. In this work, we have rare access to real-world transaction-level supply chains data, which we use to evalaute our models. We also build a new open-source simulator, SupplySim, which generates realistic supply chain data that matches the real data on key characteristics. Our simulator enables us to share data, encourage further research development on TPGs, and test models under controlled and varied settings (e.g., supply shocks, missing data). In summary, our contributions are:
@@
    Results: experiments on two real-world supply chain datasets and synthetic datasets from SupplySim, showing that our models effectively learn production functions (outperforming baselines by 6-50%) and forecast future edges (outperforming baselines by 11-62%).
While we focus on supply chains in this work, our models could be applied or easily extended to other TPGs. Our work pushes forward the frontier of graph ML and its application to real-world problems, by highlighting a new type of real-world graph and introducing new methods to handle them. [^1] Our code to run all experiments, simulator SupplySim, and synthetic datasets are available at https://github.com/snap-stanford/supply-chains/tree/tgb.
2 Learning from temporal production graphs
@@ 2.1 Problem definition
In this setting, production functions define how firms internally transform the products that they buy into products that they supply. Specifically, the function $\mathcal{F}_{p}:\mathbb{R}_{+}\rightarrow\mathbb{R}^{m}_{+}$ for product $p$ defines how much of each product is necessary to make $k$ amount of $p$ (e.g., a car requires four wheels). The set of production functions define a production graph, $G_{\textrm{prod}}$, which is a directed acyclic graph where there is an edge from products $p_{1}$ to $p_{2}$ if $p_{1}$ is required to make $p_{2}$. [^2] For simplicity, we assume in this work that there is one way to make each product, but future work may consider extensions, such as firm-specific product graphs or products that are substitutes for each other. Given the set of transactions $\mathcal{T}$, and the resulting temporal graph $G_{\textrm{txns}}$, our goals are two-fold: (1) to infer $G_{\textrm{prod}}$, which is entirely unobserved, (2) to predict future transactions, i.e., future hyperedges in $G_{\textrm{txns}}$.
@@ 2.2 Model architecture
To learn from TPGs, we introduce a new class of models that combine temporal GNNs with a novel inventory module to jointly learn production functions and predict future edges. First, we describe our inventory module, which can either operate as a stand-alone model or be attached to any GNN. Second, we describe extended versions of two popular temporal GNNs, Temporal Graph Network (Rossi et al.,, 2020) and GraphMixer (Cong et al.,, 2023), which we refer to as SC-TGN (shown in Figure 1b) and SC-GraphMixer, respectively, with SC standing for supply chain.
2.2.1 Inventory module
The basic idea of our inventory module is that it explicitly represents each firm’s inventory of products, and it adds and subtracts from the inventory based on products bought and products consumed, respectively. Let $\mathbf{x}_{i}^{(t)}\in\mathbb{R}_{+}^{m}$ represent firm $i$’s inventory at time $t$. We compute the total amount of product $p$ bought by firm $i$ at time $t$ as
@@ 2.2.1 Inventory module
$\textrm{consume}(i,p,t)$ $=\sum_{e(i,b,p_{s},t)\in\mathcal{E}}\alpha_{p_{s}p}\cdot\textrm{amt}(i,b,p_{s},t),$ (2)
@@
We take the elementwise max with 0 to ensure that the inventory stays non-negative, but we also penalize whenever consumption exceeds the current inventory amounts (see Eq. 6 below).
Attention weights.
If the inventory module is standalone, then we directly learn the pairwise weights $\alpha_{p_{1}p_{2}}$ for all product pairs $p_{1},p_{2}$. If the inventory module has access to product embeddings (e.g., from a GNN), we can use product embedding $\mathbf{z}_{p}\in\mathbb{R}^{d}$ to inform the attention weights as
@@
Inventory loss.
To train the inventory module, we introduce a special loss function. For a given firm $i$ at time $t$, its inventory loss is
$\ell_{\textrm{inv}}(i,t)$ $=\lambda_{\textrm{debt}}\sum_{p\in[m]}\max(0,\textrm{consume}(i,p,t)-\mathbf{x}_{i}^{(t)}[p])-\lambda_{\textrm{cons}}\sum_{p\in[m]}\textrm{consume}(i,p,t).$ (5)
@@
2.2.2 SC-TGN
The first GNN we explore is Temporal Graph Network (TGN) (Rossi et al.,, 2020), which is one of the most established GNNs for dynamic link prediction, outperforming other models in the Temporal Graph Benchmark (Huang et al.,, 2023). In our work, we have extended TGN to SC-TGN, by enabling it to (1) perform message passing over hypergraphs, since we represent each transaction as an edge between three nodes, (2) predict edge weights (i.e., transaction amounts) in addition to edge existence. We also modified TGN in other ways that improved performance; we document these changes in Appendix A.1. In SC-TGN, each node $i$ has a time-varying memory $\mathbf{m}_{i}^{(t)}$. Each transaction $e(s,b,p,t)$ sends three messages: one each to the supplier $s$, buyer $b$, and product $p$. At the end of each timestep, each node aggregates the messages it received and updates its memory, using a recurrent neural network (RNN). Then, to produce node embedding $\mathbf{z}_{i}^{(t)}$, we apply a GNN to the node memories, so that nodes can also learn from their neighbors’ memories (this is useful for preventing staleness, if a node has not had a transaction in a while).
2.2.3 SC-GraphMixer
We also explore GraphMixer (Cong et al.,, 2023), a recent model that showed that temporal GNNs do not always need complicated architectures, such as RNNs or self-attention (both of which are used by TGN), and strong performance can sometimes be achieved with simpler models that only rely on multi-layer perceptrons (MLPs). We similarly extend GraphMixer to SC-GraphMixer, so that it can handle hypergraphs and predict edge weight in addition to edge existence (Appendix A.2). In SC-GraphMixer, each node’s embedding $\mathbf{z}_{i}^{(t)}$ is a concatenation of its node encoding and link encoding. The node encoding is simply a sum of the node’s features and mean-pooling over its 1-hop neighbors’ features. The link encoding summarizes the recent edges that the node participated in, by applying an MLP to the concatenated time encodings and features of the recent edges.
2.2.4 Decoder
Both models, SC-TGN and SC-GraphMixer, use the same decoder architecture, and we use the same architecture (although separate decoders) for predicting edge existence and edge weight. We model these as a two-step process: first, predicting whether an edge exists; second, conditioned on the edge existing, predicting its weight. The decoder is a two-layer MLP over the concatenation of the supplier firm’s, buyer firm’s, and product’s embeddings, producing $\hat{y}\in\mathbb{R}$:
$\hat{y}(s,b,p,t)$ $=\textrm{DEC}(\mathbf{z}_{s}^{(t)},\mathbf{z}_{b}^{(t)},\mathbf{z}_{p}^{(t)})=\textrm{MLP}([\mathbf{z}_{s}^{(t)}|\mathbf{z}_{b}^{(t)}|\mathbf{z}_{p}^{(t)}]).$ (7)
@@ 2.2.4 Decoder
When the inventory module is attached, we may also allow it to inform edge prediction. To help with predicting edge existence, the inventory module penalizes impossible transactions, i.e., some $(s,b,p,t)$ where the supplier $s$ does not have enough parts to make product $p$:
Then, our model’s new prediction for edge existence becomes the sum of the original $\hat{y}$ in (7) and the penalty (8). To help with predicting edge weight, i.e., transaction amount, the inventory module computes a cap on the maximum amount of product $p$ that supplier $s$ could produce at this time:
$cap(s,b,p,t)$ $=\min_{p^{\prime}\in[m];\alpha_{pp^{\prime}}>0}\{\frac{\mathbf{x}_{s}^{(t)}[p^{\prime}]}{\alpha_{pp^{\prime}}}\}.$ (9)
Then, our new predicted amount is the minimum of the original predicted amount (7) and the cap (after applying the same log-scaling and standard scaling to the cap that we did to all amounts). Thus, as shown in Figure 1b, the inventory module and GNN help each other, with the GNN’s embeddings informing the inventory module’s attention weights, and the inventory module’s penalties and caps affecting future edge prediction.

### v3

# Yu Qin2,3, Frank Kloster2, Xi Luo2, Raj Palleti1,2, Jure Leskovec1

@@ Abstract
The global economy relies on the flow of goods over supply chain networks, with nodes as firms and edges as transactions between firms. While we may observe these external transactions, they are governed by unseen production functions, which determine how firms internally transform the input products they receive into output products that they sell. In this setting, it can be extremely valuable to infer these production functions, to improve supply chain visibility and to forecast future transactions more accurately. However, existing graph neural networks (GNNs) cannot capture these hidden relationships between nodes’ inputs and outputs. Here, we introduce a new class of models for this setting by combining temporal GNNs with a novel inventory module, which learns production functions via attention weights and a special loss function. We evaluate our models extensively on real supply chains data and data generated from our new open-source simulator, SupplySim. Our models successfully infer production functions, outperforming the strongest baseline by 6%–50% (across datasets), and forecast future transactions, outperforming the strongest baseline by 11%–62%.
@@ 1 Introduction
Supply chains form the backbone of the global economy and disruptions can have enormous consequences, costing trillions of dollars (Baumgartner, Malik, and Padhi 2020) and risking national security (The White House 2021). Thus, modeling supply chains and how they evolve, especially under shocks, is essential. Prior models of supply chains have been mainly mechanistic and struggle to fit even highly aggregated measures, such as country-level production (Inoue and Todo 2019). As a result, recent literature has called for integration of machine learning (ML) into supply chain modeling (Baryannis, Dani, and Antoniou 2019; Brintrup et al. 2020; Younis, Sundarakani, and Alsharairi 2022).
Graph neural networks (GNNs) serve as a particularly promising ML methodology, as supply chains are naturally represented as graphs, with nodes as firms and edges as transactions between firms. A few works have explored static GNNs for supply chains (Aziz et al. 2021; Kosasih and Brintrup 2021; Wasi, Islam, and Akib 2024), but these static models miss crucial dynamic aspects of supply chains, such as those driving the propagation of shocks. Thus, there is a need to develop temporal GNNs that can capture the unique dynamics and mechanisms of supply chain graphs.
@@
One key feature of these graphs is that they are governed by underlying production functions: firms receive input products (e.g., wheels) from other firms, internally transform those inputs into outputs (e.g., cars) via production functions, then sell those output products to other firms or consumers (Carvalho and Tahbaz-Salehi 2019). These production functions dictate which firms are connected to each other, as well as the timing of transactions across the network. For example, if a firm experiences a shortage in an input, its production of outputs that rely on that input will be disrupted, but it can continue to produce other outputs for some time. However, existing temporal GNNs (Huang et al. 2023) are not designed to learn these hidden production functions or to incorporate them into link prediction.
@@
The present work. Here, we are the first to develop temporal GNNs for supply chains, demonstrating their ability to learn rich, dynamic representations of firms and products. Furthermore, by introducing supply chains as a new application, we identify a challenging setting unexplored by prior GNNs: temporal graphs governed by production functions, which we term temporal production graphs (TPGs). Beyond supply chains, TPGs also appear in biological domains, such as enzymes producing metabolites within metabolic pathways, or in organizations, where teams rely on inputs from other teams to produce their outputs (e.g., engineers interfacing with product designers). TPGs introduce new challenges compared to other temporal graphs, such as the need to infer production functions. Furthermore, even standard tasks, like link prediction, may be complicated under TPGs. For example, while standard temporal GNNs might capture disruptions generally propagating across connected firms, they will miss the specific connections between each firm’s inputs and outputs and, under a shortage of inputs, not be able to precisely predict which outputs will be affected.
@@
Since existing GNNs cannot handle TPGs, we introduce a new class of GNNs designed for TPGs, focusing on two objectives: (1) learning the graph’s production functions, (2) predicting its future edges. Prior temporal GNNs have focused primarily on the second objective, but are not designed for the first. Our models support both objectives, by combining temporal GNNs with a novel inventory module, which learns production functions by explicitly representing each firm’s inventory and updating it based on attention weights that map external supply to internal consumption. Our module can be combined with any GNN; to demonstrate this, here we combine it with two popular models, Temporal Graph Network (Rossi et al. 2020) and GraphMixer (Cong et al. 2023), which we also extend in important ways.
We demonstrate the utility of our models on real and synthetic supply chains data. Through an industry collaboration, we have rare access to real transaction-level supply chains data, which allows us to evaluate our GNNs’ abilities to predict individual, future transactions. However, since we cannot release the proprietary data, we also build a new open-source simulator, SupplySim, which generates realistic supply chain data that matches the real data on key characteristics. Our simulator enables us to share data, test models under controlled and varied settings (e.g., supply shocks, missing data), and encourage future ML research on supply chains and TPGs. In summary, our contributions are:
@@
    Results: experiments on real and synthetic data, showing that our models effectively learn production functions (outperforming baselines by 6-50%) and forecast future edges (outperforming baselines by 11-62%).
Our work both contributes to the supply chains domain, as the first to develop temporal GNNs for supply chains, and advances graph ML, by introducing a new type of real-world graph and developing new methods to handle them. [^1] Our code to run all experiments, simulator SupplySim, and synthetic datasets are available at https://github.com/snap-stanford/supply-chains/tree/tgb.
Social impact.
Our work is a collaboration with Hitachi America, Ltd., a multinational conglomerate that produces a wide range of products, each involving a complex global supply chain. Our project was initiated by the company’s need for graph-based ML solutions for supply chains. By working together, we identified two ML objectives grounded in real business needs. Our first objective of learning production functions aims to improve supply chain visibility, so that firms can better understand the entire production process through which their products are produced (instead of only their immediate suppliers), identify bottlenecks, and find more efficient solutions (Aigner and Chu 1968; Coelli et al. 2005). Our second objective of forecasting future transactions supports critical industry tasks, such as demand forecasting (Seyedan and Mafakheri 2020), early detection of risks (Sheffi 2015), and inventory optimization (Vandeput 2020). Additionally, our collaboration provides access to rare, transactions-level data, which enables us to rigorously evaluate our ML models and build a realistic simulator so that, despite the data sharing constraints of this domain, we can still encourage future ML research on supply chains.
2 Related Work
Many real-world systems can be represented as temporal graphs, such as transportation systems (Jiang and Luo 2022), human mobility (Chang et al. 2023), and biological networks (Prill, Iglesias, and Levchenko 2005). While most GNNs are designed for static graphs, there has been growing interest recently in developing GNNs for temporal graphs (Huang et al. 2023; Skarding, Gabrys, and Musial 2021; Longa et al. 2023). However, to the best of our knowledge, GNNs have not yet been designed for temporal production graphs (TPGs), as described in this work. Furthermore, only a handful of works have explored GNNs for supply chains, all in static settings, such as to predict hidden links between firms (Aziz et al. 2021; Kosasih and Brintrup 2021), classify a firm’s industry (Wu, Wang, and Olson 2023), recommend suppliers (Tu et al. 2024), and predict product relations (e.g., same product group) (Wasi, Islam, and Akib 2024).
Within the supply chains domain, our work builds on prior literature that represents supply chains as networks (Fujiwara and Aoyama 2010) and models the propagation of shocks over these networks (Acemoglu et al. 2012; Zhao, Zuo, and Blackhurst 2019; Li et al. 2020; Carvalho et al. 2021). Our work is unique in two key ways: first, much of the prior work relies on theoretical models and synthetic networks, and, even among empirical studies, the data used is typically industry-level or, at best, firm-level with static relations between firms (Carvalho and Tahbaz-Salehi 2019). In contrast, we have access to transaction-level data, which reveals essential time-varying information. Second, integration of ML into supply chains has been limited, with calls for further exploration (Baryannis, Dani, and Antoniou 2019; Brintrup et al. 2020; Younis, Sundarakani, and Alsharairi 2022). Most prior models for modeling supply chains are mechanistic (Hallegatte 2008; Guan et al. 2020; Inoue and Todo 2020; Li et al. 2021) and limited in their ability to even fit aggregate counts (Inoue and Todo 2019). In contrast, by combining our inventory module with GNNs, we can accurately forecast individual transactions, while maintaining the interpretability of the inventory module.
This theme of combining deep learning with domain-specific principles also appears in physics simulation (Wang, Walters, and Yu 2021) and epidemiological forecasting (Liu et al. 2024). Our work also has connections to temporal causal discovery (Nauta, Bucur, and Seifert 2019; Löwe et al. 2022; Assaad, Devijver, and Gaussier 2022) and inferring networks from node marginals (Kumar et al. 2015; Maystre and Grossglauser 2017; Chang et al. 2024).
3 Learning on Temporal Production Graphs
@@ 3.1 Problem definition
In this setting, production functions define how firms internally transform the products that they buy into products that they supply. Specifically, the function $\mathcal{F}_{p}:\mathbb{R}_{+}\rightarrow\mathbb{R}^{m}_{+}$ for product $p$ defines how much of each product is necessary to make $k$ amount of $p$ (e.g., a car requires four wheels). The set of production functions define a production graph, $G_{\textrm{prod}}$, which is a directed acyclic graph where there is an edge from products $p_{1}$ to $p_{2}$ if $p_{1}$ is required to make $p_{2}$. [^2] For simplicity, we assume in this work that there is one way to make each product, but future work may consider extensions, such as firm-specific product graphs or substitute products. Given the set of transactions $\mathcal{T}$, and the resulting temporal graph $G_{\textrm{txns}}$, our goals are two-fold: (1) to infer $G_{\textrm{prod}}$, which is entirely unobserved, (2) to predict future transactions, i.e., future hyperedges in $G_{\textrm{txns}}$.
@@ 3.2 Model architecture
To learn from TPGs, we introduce a new class of models that combine temporal GNNs with a novel inventory module to jointly learn production functions and predict future edges. First, we describe our inventory module, which can either operate as a stand-alone model or be attached to any GNN. Second, we describe extended versions of two popular temporal GNNs, Temporal Graph Network (Rossi et al. 2020) and GraphMixer (Cong et al. 2023), which we refer to as SC-TGN (shown in Figure 1b) and SC-GraphMixer, respectively, with SC standing for supply chain.
Inventory module.
The basic idea of our inventory module is that it explicitly represents each firm’s inventory of products, and it adds and subtracts from the inventory based on products bought and consumed, respectively. Let $\mathbf{x}_{i}^{(t)}\in\mathbb{R}_{+}^{m}$ represent firm $i$’s inventory at time $t$. We compute the total amount of product $p$ bought by firm $i$ at time $t$ as
@@
$\textrm{cons}(i,p,t)$ $=\sum_{e(i,b,p_{s},t)\in\mathcal{E}}\alpha_{p_{s}p}\cdot\textrm{amt}(i,b,p_{s},t),$ (2)
@@
We take the elementwise max with 0 to ensure that the inventory stays non-negative, but we also penalize whenever consumption exceeds the current inventory amounts (6).
Attention weights. If the inventory module is standalone, then we directly learn the pairwise weights $\alpha_{p_{1}p_{2}}$ for all product pairs $p_{1},p_{2}$. If the inventory module has access to product embeddings (e.g., from a GNN), we can use product embedding $\mathbf{z}_{p}\in\mathbb{R}^{d}$ to inform the attention weights as
@@
Inventory loss. To train the inventory module, we introduce a special loss function. For a given firm $i$ at time $t$, its inventory loss is
$\ell_{\textrm{inv}}(i,t)=\lambda_{\textrm{debt}}\sum_{p\in[m]}$ $\max(0,\textrm{cons}(i,p,t)-\mathbf{x}_{i}^{(t)}[p])$ (5)
$-\lambda_{\textrm{cons}}\sum_{p\in[m]}\textrm{cons}(i,p,t).$
@@
SC-TGN.
The first GNN we explore is Temporal Graph Network (TGN) (Rossi et al. 2020), which is one of the most established GNNs for dynamic link prediction, outperforming other models in the Temporal Graph Benchmark (Huang et al. 2023). In our work, we have extended TGN to SC-TGN, by enabling it to (1) perform message passing over hypergraphs, since we represent each transaction as an edge between three nodes, (2) predict edge weights (i.e., transaction amounts) in addition to edge existence. We also modified TGN in other ways that improved performance; we document these changes in Appendix A.1. In SC-TGN, each node $i$ has a time-varying memory $\mathbf{m}_{i}^{(t)}$. Each transaction $e(s,b,p,t)$ sends three messages: one each to the supplier $s$, buyer $b$, and product $p$. At the end of each timestep, each node aggregates the messages it received and updates its memory, using a recurrent neural network (RNN). Then, to produce node embedding $\mathbf{z}_{i}^{(t)}$, we apply a GNN to the node memories, so that nodes can also learn from their neighbors’ memories (this is useful for preventing staleness, if a node has not had a transaction in a while).
SC-GraphMixer.
We also explore GraphMixer (Cong et al. 2023), a recent model that showed that temporal GNNs do not always need complicated architectures, such as RNNs or self-attention (both of which are used by TGN), and strong performance can sometimes be achieved with simpler models that only rely on multi-layer perceptrons (MLPs). We similarly extend GraphMixer to SC-GraphMixer, so that it can handle hypergraphs and predict edge weight in addition to edge existence (Appendix A.2). In SC-GraphMixer, each node’s embedding $\mathbf{z}_{i}^{(t)}$ is a concatenation of its node encoding and link encoding. The node encoding is simply a sum of the node’s features and mean-pooling over its 1-hop neighbors’ features. The link encoding summarizes the recent edges that the node participated in, by applying an MLP to the time encodings and features of the recent edges.
Decoder.
Both models, SC-TGN and SC-GraphMixer, use the same decoder architecture, and we use the same architecture (although separate decoders) for predicting edge existence and weight. We model these as a two-step process: first, predicting whether an edge exists; second, conditioned on the edge existing, predicting its weight. The decoder is a two-layer MLP over the concatenated supplier firm’s, buyer firm’s, and product’s embeddings, producing $\hat{y}\in\mathbb{R}$:
$\hat{y}(s,b,p,t)$ $=\textrm{MLP}([\mathbf{z}_{s}^{(t)}|\mathbf{z}_{b}^{(t)}|\mathbf{z}_{p}^{(t)}]).$ (7)
@@
When the inventory module is attached, we may also allow it to inform edge prediction. For edge existence, the inventory module penalizes impossible transactions, i.e., some $(s,b,p,t)$ where supplier $s$ does not have the parts required to make product $p$ in its inventory at time $t$, and subtracts the penalty (14) from the model’s original $\hat{y}$ (7). For edge weight, the inventory module computes the maximum amount of product $p$ that supplier $s$ could produce, based on its current inventory, and caps $\hat{y}$ based on the computed maximum (15). Thus, as shown in Figure 1b, the inventory module and GNN help each other, with the GNN’s embeddings informing the inventory module’s attention weights, and the inventory module’s penalties and caps affecting future edge prediction.